OpenStudy (anonymous):

|-17| + 32 = 80y - |-31| - 79y =57??

OpenStudy (anonymous):

too many equal signs

OpenStudy (hoblos):

|-17| + 32 = 80y - |-31| - 79y 17 +32 = y - 31 y= 80

Directrix (directrix):

y = 80 --> I saw only one = sign.

OpenStudy (anonymous):

maybe last line meant y = 57, nvm

OpenStudy (anonymous):

BUT LOOOOOK

Directrix (directrix):

9_9 I see two = now.

OpenStudy (anonymous):

ahh \[3^2\] not 32

Directrix (directrix):

57 now with only one =

OpenStudy (anonymous):

57 is right

OpenStudy (hoblos):

17 +9 = y - 31 y= 57 now you are right

Directrix (directrix):

That would be A) 57